Refer to the table at right. The four-firm concentration ratio is Firm Annual Sales OA. 72.5 percent. O B. 75 percent. O C. 59.2 percent. O D. 85.8 percent. 800 75 50 40 800 1,500 1,050 90 75 300
Answer #1

Ans: 72.5 percent

Explanation:

Total Sales = 1000 + 800 + 220 + 75 + 50 + 40 + 800 + 1500 + 1050 + 90 + 75 + 300 = $6,000

Share of four largest firm = 1500 + 1050 + 1000 + 800 = $4,375

Four Firm concentration ratio = (4,375 / 6,000) * 100 = 72.5 percent